MrQuestion:

Question 4(Multiple Choice Worth 5 points) (06.06 LC) Which expression is equal to a + (b + c)? (a + b) + c (a + b) ⋅ c a + bc b + ac Question 5(Multiple Choice Worth 5 points) (06.03 LC) What is the constant term in the expression 8xy + 9x2 y − 5x − 2? −5 −2 8 9 Question 6(Multiple Choice Worth 5 points) (06.07 MC) Four expressions are shown below: 2(4x + 2) 2(3x + 3) 6x + 6 8x + 4 Which two expressions are equivalent to 2(3x + 2 + x)? A and D B and D A and C C and B Question 7(Multiple Choice Worth 5 points) (06.01 MC) What is the value of the expression shown below? 6 + (9 − 1) 2 ÷ 4 ⋅ 1 over 2 whole to the power of 4 7 9 18 22 Question 8 (Fill-In-The-Blank Worth 5 points) (06.04 LC) If s = 1 over 3 unit and A = 18s2, what is the value of A, in square units? ____ square unit(s). (Input whole number only.) Numerical Answers Expected! Answer for Blank 1: Question 9(Multiple Choice Worth 5 points) (06.02 LC) Which expression represents the product of 6 and y? 6y 6 + y 6 − y 6 over y Question 10(Multiple Choice Worth 5 points) (06.03 MC) Which of the following best describes the expression 12(x + 3)? The product of a constant factor 12 and a 2-term factor x + 3 The sum of a constant factor 12 and a 2-term factor x + 3 The sum of constant factors 12 and x + 3 The product of constant factors 12 and x + 3 Question 11(Multiple Choice Worth 5 points) (06.07 LC) Which expression is equivalent to 5x + 5x? 25x 10 + 2x 25x2 10x Question 12(Multiple Choice Worth 5 points) (06.02 LC) What best describes the expression y over 4? Some number divided by 4 Some number times 4 4 more than some number 4 less than some number

supie:

Any ideas on question 1?

MrQuestion:

A times B + C But thats all i got

supie:

@jhonyy9 can you explain this pls ? ty

MrQuestion:

I still got nothing

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie @jhonyy9 can you explain this pls ? ty \(\color{#0cbb34}{\text{End of Quote}}\) the user is offline now

supie:

oh

MrQuestion:

Im bck

jhonyy9:

on what part you need help ?

MrQuestion:

Part 8

jhonyy9:

what is this ? from the top the 8th ?

jhonyy9:

ok this one If s = 1 over 3 unit and A = 18s2, what is the value of A, in square units? ____ square unit(s). (Input whole number only.) Numerical Answers Expected!

jhonyy9:

yes ?

MrQuestion:

No the input whole nimbers one

MrQuestion:

Thats the one i need help on

jhonyy9:

@supie what is this ? pls

jhonyy9:

this is the problem - yes ?

MrQuestion:

I need help on Question 8 (Fill-In-The-Blank Worth 5 points) (06.04 LC) If s = 1 over 3 unit and A = 18s2, what is the value of A, in square units? ____ square unit(s). (Input whole number only.) Numerical Answers Expected! Answer for Blank 1:

supie:

\(s=\frac13\) and \(A=18s^2\) ?

jhonyy9:

If s = 1 over 3 unit and A = 18s2, what is the value of A, in square units? ____ square unit(s). (Input whole number only.) Numerical Answers Expected! \[s = \frac{ 1 }{ 3 } and A = 18s^2 = 18(\frac{ 1 }{ 3 })^2\] A = ?

jhonyy9:

the user go away

supie:

They'll come back and see.

jhonyy9:

@supie can you solve it now ?

supie:

What do I solve...?

jhonyy9:

solve this for A

supie:

\(A = 18s^2 = 18(\frac{ 1 }{ 3 })^2\) this?

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 If s = 1 over 3 unit and A = 18s2, what is the value of A, in square units? ____ square unit(s). (Input whole number only.) Numerical Answers Expected! \[s = \frac{ 1 }{ 3 } and A = 18s^2 = 18(\frac{ 1 }{ 3 })^2\] A = ? \(\color{#0cbb34}{\text{End of Quote}}\) this equation

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie \(A = 18s^2 = 18(\frac{ 1 }{ 3 })^2\) this? \(\color{#0cbb34}{\text{End of Quote}}\) yes

supie:

\(18\frac19\)?

jhonyy9:

\[18 *(\frac{ 1 }{ 9 }) = ?\]

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie \(18\frac19\)? \(\color{#0cbb34}{\text{End of Quote}}\) this maybe 18 integer and fraction what is different

jhonyy9:

how you multiplie integer by a fraction ?

jhonyy9:

\[\frac{ 18 }{ 1 }×\frac{ 1 }{ 9 } = ?\]

supie:

Yes so \(\large 18 \times\frac19=\frac{18}{1} \times \frac19 \) so then I just multiply straight across, \(\frac{18}{1} \times \frac19=\frac{18}{9}=2 \)

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie Yes so \(\large 18 \times\frac19=\frac{18}{1} \times \frac19 \) so then I just multiply straight across, \(\frac{18}{1} \times \frac19=\frac{18}{9}=2 \) \(\color{#0cbb34}{\text{End of Quote}}\) explained perfect - ty

supie:

np thanx